A certain test preparation course is designed to help students improve their scores on the LSAT exam. A mock exam is given at the beginning and end of the course to determine the effectiveness of the course. The following measurements are the net change in 3 students' scores on the exam after completing the course:

24,18,11

Using these data, construct a 95%confidence interval for the average net change in a student's score after completing the course. Assume the population is approximately normal.
Answer #1

sample mean, xbar = 17.667
sample standard deviation, s = 6.5064
sample size, n = 3
degrees of freedom, df = n - 1 = 2

Given CI level is 95%, hence α = 1 - 0.95 = 0.05
α/2 = 0.05/2 = 0.025, tc = t(α/2, df) = 4.303


ME = tc * s/sqrt(n)
ME = 4.303 * 6.5064/sqrt(3)
ME = 16.164

CI = (xbar - tc * s/sqrt(n) , xbar + tc * s/sqrt(n))
CI = (17.667 - 4.303 * 6.5064/sqrt(3) , 17.667 + 4.303 * 6.5064/sqrt(3))
CI = (1.50 , 33.83)